OpenStudy (anonymous):

Adding Two Dot Products How does r'(t)·r(t) + r(t)·r'(t) = 2r'(t)·r(t) ?

OpenStudy (phi):

the dot product commutes. let A= (a,b) and B= (x,y) A dot B = a*x + b*y B dot A = x*a + y*b but x*a = a*x and y*b= b*y B dot A matches A dot B

OpenStudy (phi):

that means A \( \cdot \) B + B ⋅ A = A ⋅ B + A ⋅ B = 2 A ⋅ B